OpenStudy (anonymous):

verify the identity cosx/1-cscx=cscx-1/cotx

OpenStudy (anonymous):

I know that I have to add sine to the equation... but thats about it

OpenStudy (mertsj):

No. To prove an identity, you cannot work across the equal sign

OpenStudy (ash2326):

\[\frac{\cos x}{1-cosec\ x}\] Multiply numerator and denominator by (1+cosec x ) \[\frac{\cos x}{1-cose\ x}\times \frac{1+cosec\ x}{1+cosec\ x}\] \[\frac{\cos x \times (1+cosec\ x)}{1-cosec^2 x}\] We know that \[1-cosec^2 x=-cot^2 x\] \[\frac{\cos x \times (1+cosec\ x)}{-\cot^2 x}\] Can you try from here?

OpenStudy (anonymous):

okay let me try from here but what about the other side?...like I don't exactly get what I am trying to accomplish?

OpenStudy (ash2326):

Only work on one side, I'm sure you'll get the same expression as right side

OpenStudy (anonymous):

okay wait.. does cos become 1/secx?

OpenStudy (anonymous):

cosx I mean

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

Yes

OpenStudy (anonymous):

okay so then I will have \[1/secx*(1+cosecx) / (-\cot^2x)\]

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

I honestly have no clue what to do after...

OpenStudy (ash2326):

\[\frac{\cos x \times (1+cosec\ x)}{-\cot^2 x}\] \[\frac{\cos x \times (1+cosec\ x)}{-\cot x\times \frac{\cos x}{\sin x}}\] Now we get \[\frac{\sin x \times (1+cosec\ x)}{-\cot x}\] \[\frac{\sin x+ 1}{-\cot x}\] Do you understand till here?

OpenStudy (anonymous):

the only thing I don't get is where did cosec x go?

OpenStudy (ash2326):

sin x =1/cosec x

OpenStudy (anonymous):

oh so it cancels out?

OpenStudy (ash2326):

yes

OpenStudy (anonymous):

okay so I get sinx+1/-cotx = -cotx/sinx+1?

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

I could see that you have put the identity wrong, if you put x=45 degrees. You can verify that left side is not equal to right side \[\frac{\frac{1}{\sqrt 2}}{1-\sqrt 2}\ne\frac{\sqrt 2-1}{1}\] Are you sure that the question is correct?

OpenStudy (anonymous):

verify the identity cosx/1+cscx=cscx-1/cotx its positive on the left side.. >.< sorry

OpenStudy (ash2326):

So we have to multiply the left side by (1-cosec x ) Now proceed the way I did.